OpenStudy (anonymous):

The equation is : px^2+qx+r=0 and 'a' is the repeated root. f(x)=px^2+qx+r=0 f(a)=p(x-a)(x-a) HOW DID THEY ARRIVE AT THIS?

OpenStudy (anonymous):

my dear then notice in the question that a is repeated root so u should make two equation and compare them with each other. coefficient of x with same degree should be equal . consider what you have . f(x)=px^2+qx+r=0 f(a)=p(x-a)(x-a) first equation is : f(x)=px^2+qx+r=0 and the second one is : f(a)=p(x-a)(x-a) =px^2+2ax+a^2=0 we suppose to compare coefficient of x^2 , they should be equal and they are . for the second part compare q with 2a so we find out that q is equal with 2a :q=2a and with the same way we find that :a^2=r

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome dear